cos (alpha+beta)= 0 then sin(alpha + 2beta ) = a. sin alpha b. cos alpha c. sin beta d. cos 2beta ​I want answer with full solution

cos(alpha+beta)= 0 then sin(alpha + 2beta) =
a. sin alpha       b. cos alpha      c. sin beta      d. cos 2beta

​I want answer with full solution

Grade:11

2 Answers

Shubham Kamleshbhai Patel
8 years ago
This is an easy question.
Shubham Kamleshbhai Patel
8 years ago
cos (α+β) = 0
 
So let (α+β) = θ.
Then cos θ = 0 => θ = 90°
 
Thus (α+β) = 90°
So, β = 90° – α -------------------------- (1)
 
Substituting (1) in sin (α+2β),
=> sin [α + 2(90° – α)]
=> sin(180° – α)
=> sin α
SO THE CORRECT OPTION IS A.